OpenStudy (anonymous):

tanA + SecA - 1/tanA - SecA +1 = 1+ SinA/CosA Solve using trigonometrical functions and componendo dividendo......solve fast....plz its urgent....

OpenStudy (anonymous):

\[\frac{ \tan A+\sec A-1 }{ \tan A-\sec A+1 } = \frac{ 1+\sin A }{ \cos A }\]

OpenStudy (anonymous):

Prove this....

OpenStudy (anonymous):

Prove that the Left hand side is equal to the right hand side.... Using functions, ratios , identities and componendo - dividendo.......Solve fast its really urgent....

OpenStudy (anonymous):

@Loser66 Alright.....no problem....but i hope not more than 30 mins.....

OpenStudy (loser66):

@amistre64

OpenStudy (loser66):

sorry friend, I am useless, cannot help. I asked someone for help. You have to wait for them

OpenStudy (anonymous):

@Loser66 alright will wait...

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

try to convert all in sin and cos :) ans solve the simple maths

OpenStudy (anonymous):

i did that i got SinA +1-cosA/ SinA-1+Cos A plz tell me how to apply the componendo dividendo here....@chmvijay

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

wait let me try wt is sec = i /cos A is it and tan A= sin A / cos a right

OpenStudy (anonymous):

yup @chmvijay

OpenStudy (chmvijay):

yaa ur right did you check the question once again

OpenStudy (loser66):

Can I go from RHS? if I can, I think I got the answer

OpenStudy (anonymous):

ya u can @Loser66

OpenStudy (chmvijay):

yaaa try @Loser66 you may good idea :)

OpenStudy (loser66):

ok

OpenStudy (loser66):